OpenStudy (anonymous):

PLEASE HELP ASAP WILL GIVE MEDAL TO FRST PERSON WHO ANSWERS!!!! A recipe calls for 10 cups of oatmeal. Benjamin wants to make 1/5 fewer servings than the recipe makes. Which expression represents how many cups of oatmeal Benjamin should use? A: B: C: D:

OpenStudy (anonymous):

A: 0.8 X 10 B: 1.8 X 10 C: 0.2 X 10 D: 1.2 X 10

sammixboo (sammixboo):

Well can you telll me what is 1/5 first :)

OpenStudy (anonymous):

thats what i need help with -.-

sammixboo (sammixboo):

Aaah.. 1/5 = 2/10 Do you know what \(\rm \frac{2}{10}\) is in decimal form?

OpenStudy (anonymous):

so D? thnx

sammixboo (sammixboo):

NO no no :) We are still solving this! Do you know what \(\rm \frac{2}{10}\) is in decimal form?

OpenStudy (anonymous):

2.1

sammixboo (sammixboo):

No it is 0.2 :) \(\rm \frac{2}{10}\) = 0.2

sammixboo (sammixboo):

And 1/5 = 0.2 So now we know that 1/5 = 0.2

OpenStudy (anonymous):

oh...yea im bad at these :/

sammixboo (sammixboo):

So to solve this problem we must do 1/5 x 10 And remember 0.2 = 1/5 SO 0.2 x 10 = 1/5 x 10

OpenStudy (anonymous):

so c?

sammixboo (sammixboo):

Yeppers!

OpenStudy (anonymous):

>w< thnx

OpenStudy (anonymous):

It's actually A. 0.8 x 10

OpenStudy (anonymous):

thanks :)